OpenStudy (anonymous):

(5-x)(3+x)=2 Find a, b and c

OpenStudy (anonymous):

What a b c?

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

15+2x-x^2 a-x^2 b-2x c-15

OpenStudy (anonymous):

Just foil it out

OpenStudy (anonymous):

formula?

sam (.sam.):

@DylanH you forgot the 2

OpenStudy (anonymous):

what 2?

mathslover (mathslover):

\[\huge{(5-x)(3+x)=2}\] \[\huge{5(3)+5x-3x+x^2=2}\] \[\huge{15+2x+x^2=2}\] The general form of a quadratic equation is \(\large{ax^2+bx+c=0}\) compare both equations : can u do this now ?

mathslover (mathslover):

Note u can also write \(\huge{15+2x+x^2=2}\) as : \[\huge{x^2+2x+15=2}\]

OpenStudy (anonymous):

ohhh my bad

OpenStudy (anonymous):

ok tnx

OpenStudy (anonymous):

what is c pls. i knew a=1 b= 2 c=?

OpenStudy (anonymous):

@mathslover PLS ANSWER

OpenStudy (anonymous):

13..

mathslover (mathslover):

a : 1 , b: 2 and c : 15

OpenStudy (anonymous):

isnt it 13? cause you will transpose 2 to the left side 15-2 =13

mathslover (mathslover):

oh yes sorry u r right

OpenStudy (anonymous):

@mathslover pls answer may last question (2/x)^2-4=7 find a b and c again pls

mathslover (mathslover):

\[\huge{\frac{4}{x^2}-4=7}\] \[\huge{\frac{4}{x^2}-11=0}\] \[\huge{\frac{4-11x^2}{x^2}=0}\] \[\huge{4-11x^2=0}\] \[\huge{11x^2-4+0=0}\] a = 11 , b = -4 , c =0 :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if the hat cost 9.53 and the sale taxes was7% how much did rosa pay for the hat ?

OpenStudy (anonymous):

$10.20

mathslover (mathslover):

@lanceglojames first of all welcome to openstudy I request you to post the questions as there in the bar of You have an open question : ask ... dont post in the other student's question .thanks!

OpenStudy (anonymous):

25 childern will attend the birthday party . if each child gets 15 party favors and there are 75 favors in each box . how many boxs should raymond buy ?